# Open the Eyes of MPNN: Vision Enhances MPNN in Link Prediction

### v2

@@ Appendix A Supplementary Algorithms
This section provides the supplementary algorithms for both GVN and E-GVN.
@@ A.2 Algorithm for Efficient GVN (E-GVN)
Algorithm 3 The Efficient GVN (E-GVN) Framework
@@ B.1 Time Complexity Analysis for GVN
    Complexity of generating visual images for the target links is $O(l)$
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(l(SF^{\prime}+F^{\prime 2}))$
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(l(F^{\prime}+S))$
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(lS^{2})$
@@ B.2 Time Complexity Analysis for E-GVN
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+SF+F^{2}))$
@@
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+F+S))$
@@
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+F^{2}))$
@@ Appendix C Dataset Statistics
Table 11: Statistics of dataset.
@@ Appendix D Link Prediction Experimental Setups
In link prediction, links play dual roles: serving as supervision and acting as message-passing paths. Following the standard practice in link prediction, training links fulfill both supervision labels and message-passing paths. In terms of supervision, the training, validation, and testing links are mutually exclusive. For message passing, we follow the common setting where the validation links in ogbl-collab additionally function as message-passing paths during test time.
@@ Appendix E Substructure Counting Experimental Details
Models We compare GCN (Kipf & Welling, 2016), VSF+GCN, SAGE (Hamilton et al., 2017), and VSF+SAGE. To predict the substructure count, we append a 3-layer MLP decoder after GCN or SAGE. For VSF+ “‘GCN and VSF+SAGE, we concatenate the VSFs with the node representations produced by the MPNN, which are then used as inputs to the MLP decoder. All models are trained for 100 epochs using the Adam optimizer (Kingma, 2014), with learning rates searched from $\{1,0.1,0.05,0.01\}$. The depths of GCN and SAGE are searched from $\{2,3,4,5\}$, and the hidden dimensions are searched from $\{8,32,128,256\}$. We select the best model based on the lowest MSE loss on the validation set to generate the results.
Appendix F Evaluation on other Metrics
We test our model in different metrics. The results are shown in Table 12. In total, E-GVN achieves 39 best scores (in bold), GVN achieves 7 best scores, and our strongest baseline NCNC achieves 3 best scores. Therefore, GVNNCNC and E-GVNNCNC still consistently outperform NCNC in more comprehensive metrics.
Table 12: Models’ performance with various metrics. NCNC is our strongest baseline.
Cora Citeseer Pubmed Collab PPA Citation2 DDI
hit@3 GVNNCNC $26.66_{\pm 5.96}$ $59.97_{\pm 6.21}$ $\mathbf{32.23_{\pm 5.69}}$
hit@10 GVNNCNC $\mathbf{58.83_{\pm 5.29}}$ $75.28_{\pm 3.03}$ $40.34_{\pm 2.28}$
hit@20 GVNNCNC $\mathbf{70.01_{\pm 4.44}}$ $81.11_{\pm 1.30}$ $53.33_{\pm 2.67}$
hit@50 GVNNCNC $82.06_{\pm 1.94}$ $88.88_{\pm 0.98}$ $\mathbf{71.66_{\pm 2.75}}$
hit@100 GVNNCNC $90.70_{\pm 0.56}$ $94.12_{\pm 0.58}$ $82.17_{\pm 0.77}$
mrr GVNNCNC $\mathbf{24.66_{\pm 4.51}}$ $62.74_{\pm 6.63}$ $26.32_{\pm 6.67}$
Appendix G Graph Visualizer
@@
Appendix H Style Consistency Ablation Experimental Details
@@
The consistent style setting is achieved by using the default image style configuration as described in Appendix G. This involves maintaining uniformity in visual elements across all subgraph images, ensuring that each image adheres to the same stylistic parameters.
@@
Appendix I Impact Study for Image Styles
I.1 Impacts of Node Colors
@@
I.2 Impacts of Node Shapes
@@
I.3 Impacts of Node Labeling Strategies
I.3.1 Sensitive Study of Node Labeling Strategies
@@
In this experiment, we study different ways to label the nodes in the image: (i) “No-label”, which shows the nodes without any labels; (ii) “Re-label”, which maps all the nodes in the current subgraph to new labels starting from zero; (iii) “Unique”, which labels the nodes with unique global indices. Example images for these visualization strategies are shown in Appendix I.3.2. Table 16 shows the HR@100 performance of GVNNCNC and E-GVNNCNC with these different labeling strategies on Cora and Citeseer. As can be seen, “no-label” performs best, indicating that purely using the structural information is preferred.
I.3.2 Illustrations for Different Labeling strategies

### v3

@@ Appendix A Supplementary Algorithms
This section provides the algorithms for GVN and E-GVN.
@@ A.2 Algorithm for Efficient GVN (E-GVN)
Algorithm 3 The Efficient GVN (E-GVN) Framework.
@@ B.1 Time Complexity Analysis for GVN
    Complexity of generating visual images for the target links is $O(l)$.
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(l(SF^{\prime}+F^{\prime 2})).$
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(l(F^{\prime}+S)).$
@@
    $O(\text{MPNN})+l\cdot O(\text{VE})+O(lS^{2}).$
@@ B.2 Time Complexity Analysis for E-GVN
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+SF+F^{2})).$
@@
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+F+S)).$
@@
    $O(\text{MPNN})+n\cdot O(\text{VE})+O(n(S^{2}+F^{2})).$
@@ Appendix C Dataset Statistics
Table 12: Statistics of datasets.
@@ Appendix D Link Prediction Experimental Setups
In link prediction, links play dual roles: serving as supervision and acting as message-passing paths. Following the standard practice in link prediction, training links fulfill both supervision labels and message-passing paths. In terms of supervision, the training, validation, and testing links are mutually exclusive. For message passing, we follow the common setting (Li et al., 2024) where the validation links in ogbl-collab additionally function as message-passing paths during test time.
@@ Appendix E Substructure Counting Experimental Details
Models We compare GCN (Kipf & Welling, 2017), VSF+GCN, SAGE (Hamilton et al., 2017), and VSF+SAGE. To predict the substructure count, we append a 3-layer MLP decoder after GCN or SAGE. For VSF+ “‘GCN and VSF+SAGE, we concatenate the VSFs with the node representations produced by the MPNN, which are then used as inputs to the MLP decoder. All models are trained for 100 epochs using the Adam optimizer (Kingma, 2014), with learning rates searched from $\{1,0.1,0.05,0.01\}$. The depths of GCN and SAGE are searched from $\{2,3,4,5\}$, and the hidden dimensions are searched from $\{8,32,128,256\}$. We select the best model based on the lowest MSE loss on the validation set to generate the results.
Appendix F Graph Visualizer
@@
Appendix G Style Consistency Ablation Experimental Details
@@
The consistent style setting is achieved by using the default image style configuration as described in Appendix F. This involves maintaining uniformity in visual elements across all subgraph images, ensuring that each image adheres to the same stylistic parameters.
@@
Appendix H Impact Study for Image Styles
H.1 Impacts of Node Colors
@@
H.2 Impacts of Node Shapes
@@
H.3 Impacts of Node Labeling Strategies
H.3.1 Sensitive Study of Node Labeling Strategies
@@
In this experiment, we study different ways to label the nodes in the image: (i) “No-label”, which shows the nodes without any labels; (ii) “Re-label”, which maps all the nodes in the current subgraph to new labels starting from zero; (iii) “Unique”, which labels the nodes with unique global indices. Example images for these visualization strategies are shown in Appendix H.3.2. Table 16 shows the HR@100 performance of GVNNCNC and E-GVNNCNC with these different labeling strategies on Cora and Citeseer. As can be seen, “no-label” performs best, indicating that purely using the structural information is preferred.
H.3.2 Illustrations for Different Labeling strategies